OpenStudy (anonymous):

Is my answer correct? Q: evaluate cos(15). Answer: Cos(45-30) = cos(45)cos(30) + sin(45)sin(30) =(1/Sq root of 2)(Sq Root of 3/2)+(1/Sq root of 2)(1/2) = (1+sq root of 3/2 sq root of 2)+(1/2 sq root of 2) =sq root of 6+ Sq root of 2/4

OpenStudy (anonymous):

first line is right, then i am confused

OpenStudy (anonymous):

oh no first line is not right, should be plus not minus

OpenStudy (anonymous):

\[Cos(45-30) = cos(45)cos(30) +sin(45)sin(30)\]

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{2}\times \frac{\sqrt{3}}{2}+\frac{\sqrt{2}}{2}\times \frac{1}{2}\]